# Force Fields for Deep Eutectic Mixtures: Application to Structure and 2D-Infrared Spectroscopy

### v1

@@ 4 Discussion and Conclusion
applications66 and to further readjust
parameters.57, 63 The adjustment is usually
to scale partial charges62 and in some cases
@@
diffusivities, viscosities or the densities.63 In terms
@@
m2/s).61, 57, 67 Thus the
determination of the values typically require a set of ns simulations (compared with 5 ns for water67)
@@
symmetry adapted perturbation theory (SAPT)61 which
@@
CL&Pol.58 Recent fine-tuning efforts used structural
data.59 The present status of computational approaches
@@
hydration free energies as for M0 and M1.
@@
previously,10 is affected by intermolecular
@@ 5. For model M0 the angular dependence of these
parametrization10 do not differ appreciably
(compare with Figure 5 in Ref.10)
@@
mPas at 303 K.69 Simulations with M1 and
@@
with those from simulations.70 As was already
@@
the TIP4P water model.27 Again, hydrated
@@
Figure 9. It is found that the two water models yield
similarly accurate representations of the intermolecular
interactions. With this new parametrization, simulations using the
TIP4P model can be carried out to determine possible changes for the
structure and spectroscopy when using a water model different from
TIP3P. This, however, is beyond the scope of the present work.
The present work combines state-of-the art electrostatic models with
established empirical energy functions to correctly describe
thermodynamic properties of components used in eutectic mixtures
consisting of KSCN, acetamide and water. The refined models are then
used to characterize structural data (radial distribution functions),
measured spectroscopic properties from 2-dimensional infrared
spectroscopy, and thermodynamic properties (viscosity). It is found
that electrostatic models based on distributed charges (M1 and
M2) perform better than a model using multipoles (M0) to
reproduce the magnitudes of the FFCF decay times for different water
contents as seen in experiments. For the radial distribution functions
it should be noted that they are not direct experimental observables
but rather fitted to reproduce neutron scattering data. Given the
performance of the models, in particular M1 and M2 on the
measured thermodynamic and spectroscopic data it is anticipated that
the computed atom-atom radial distributions functions from simulations
are even more reliable than those determined from experiment.
In summary, the parametrization strategies pursued in the present work
successfully combined electronic structure methods, comparison with
thermodynamic reference data and application to measured structural
and spectroscopic properties. Adjustment of both, electrostatic
contributions and LJ-parameters was found to be important. The present
work suggests that distributed charge models are preferable over
multipole-based representations of the electrostatics for the system
compositions considered here. Application of the procedures outlined
here to other, related mixtures and additional observables will
provide further insight into generalizability and scope for
improvements of this approach.

### v2

@@ 4 Discussion and Conclusion
applications78 and to further readjust
parameters.31, 37 The adjustment is usually
to scale partial charges36 and in some cases
@@
diffusivities, viscosities or the densities.37 In terms
@@
m2/s).35, 31, 79 Thus the
determination of self-diffusion coefficients typically requires a set of ns simulations (compared with 5 ns for water79)
@@
symmetry adapted perturbation theory (SAPT)35 which
@@
CL&Pol.32 Recent fine-tuning efforts used structural
data.33 The present status of computational approaches
@@
hydration free energies as for M0 and M1. This
strategy follows common practices in empirical energy function
development in that the electrostatic parameters usually originate
from high-level quantum chemical calculations whereas the
van-der-Waals parameters are adjusted with respect to, for example, interaction
energies and hydration free energies.80
@@
previously,9 is affected by intermolecular
@@ 5. For model M0 the angular dependence of these
parametrization9 do not differ appreciably
(compare with Figure 5 in Ref.9)
@@
mPas at 303 K.82 Simulations with M1 and
@@
with those from simulations.83 As was already
@@
the TIP4P water model.46 Again, hydrated
@@
Figure 9 and Table S4. It is
found that the two water models yield similarly accurate
representations of the intermolecular interactions. With this new
parametrization, simulations using the TIP4P model can be carried out
to determine possible changes for the structure and spectroscopy when
using a water model different from TIP3P. This, however, is beyond the
scope of the present work. Another future extension of the present work is to use kernel-based MDCM (kMDCM) to capture conformationally dependent changes in the ESP in particular for molecules larger than triatomics.84
In summary, the parametrization strategies pursued in the present
work successfully combined electronic structure methods, comparison
with thermodynamic reference data and application to measured
structural and spectroscopic properties. Adjustment of both,
electrostatic contributions and LJ-parameters was found to be
important. The present work suggests that distributed charge models
are preferable over multipole-based representations of the
electrostatics for the system compositions considered
here. Application of the procedures outlined here to other, related
mixtures and additional observables will provide further insight
into generalizability and scope for improvements of this
approach.